OpenStudy (anonymous):

solve 8m^2 + 20m= 12 for m by factoring

OpenStudy (anonymous):

4(2m^2+5m)=12

Nnesha (nnesha):

subtract 12 both side first

OpenStudy (anonymous):

Thanks

Nnesha (nnesha):

we are not done .............

Nnesha (nnesha):

:)

OpenStudy (idealist10):

8m^2+20m=12 factor out the 4 2m^2+5m=3 move 3 to the other side 2m^2+5m-3=0 factor out the trinomial, find 2 numbers that when they multiply, you get -3 and when they add up, you get 5. (2m-1)(m+3)=0 zero property 2m-1=0 and m+3=0 2m=0+1=1 m=1/2 ----------------- m=0-3=-3 Answer: 1/2, -3.

Nnesha (nnesha):

 that's what you should get after subtracting 12 both side can you find GCF now ??